### v1

Abstract
In soccer video analysis, player detection is essential for identifying key events and reconstructing tactical positions. The presence of numerous players and frequent occlusions, combined with copyright restrictions, severely restricts the availability of datasets, leaving limited options such as SoccerNet-Tracking and SportsMOT. These datasets suffer from a lack of diversity, which hinders algorithms from adapting effectively to varied soccer video contexts. To address these challenges, we developed SoccerSynth-Detection, the first synthetic dataset designed for the detection of synthetic soccer players. It includes a broad range of random lighting and textures, as well as simulated camera motion blur. We validated its efficacy using the object detection model (Yolov8n) against real-world datasets (SoccerNet-Tracking and SportsMoT). In transfer tests, it matched the performance of real datasets and significantly outperformed them in images with motion blur; in pre-training tests, it demonstrated its efficacy as a pre-training dataset, significantly enhancing the algorithm’s overall performance. Our work demonstrates the potential of synthetic datasets to replace real datasets for algorithm training in the field of soccer video analysis.
keywords:
synthetic datasets, object detection, soccer
1 Introduction
Player detection serves as a foundational task in sports video processing, directly influencing subsequent tasks such as tracking [1, 2], action recognition [3, 4, 5, 6, 7], action prediction [8], and game state reconstruction [9, 10]. Traditional detection methods such as the Viola-Jones Detector [11], Histograms of Oriented Gradients [12], and Discriminatively Trained Part-Based Models [13] rely on manually designed features. In contrast, deep learning-based algorithms like YOLO (You Only Look Once) [14, 15, 16, 17, 18, 19, 20, 21] enhance detection speed through a unique single-shot approach, offering real-time, robust detection and significant evolutionary improvements. Faster R-CNN [22] also contributes to the field with its advanced capabilities.
Despite significant advancements in detection algorithms, the application to sports scenarios encounters challenging due to complexities in annotating datasets, including frequent occlusions, dynamic backgrounds, lighting variations, motion-induced blurring, and copyright restrictions. Among various sports, soccer stands out with its large number of players, which exacerbates occlusions and complicates tracking. There are several datasets for the detection of team sports players, such as SoccerNet-Tracking [23], a tracking dataset from live soccer match recordings; SoccerTrack [2] and TeamTrack [24], using fisheye cameras and drones to cover various sports; SoccerDB [25], a target detection dataset from TV broadcasts; 3x3 basketball datasets [26, 1]; and SportsMot [27], designed for multi-sport player tracking. However, the overall availability of datasets for the detection of soccer players remains limited.
Employing simulators to generate synthetic datasets has emerged as a promising solution to address dataset scarcity, gaining significant attention in recent years for enabling large-scale, low-cost, and precise label generation across domains such as pedestrian detection [28, 29, 30], autonomous driving [31], and indoor scene understanding [32, 33]. In the field of soccer, several studies have attempted to create synthetic datasets. For instance, the SoccerNet-Depth dataset [34] employs NVIDIA Nsight Graphics to extract depth maps of players from eFootball games. The RADepthNet study [35] produced a soccer depth map dataset known as the FIFADataset, using the open-source tool RenderDoc. Web-based tools are utilized to generate soccer field datasets for keypoint extraction [36]. Furthermore, Kaneko et al. [37] implemented the Google Research Football simulator [38] to train algorithms aimed at predicting pass movements in soccer. Additionally, Qin et al. [39] outline a methodology that employs a soccer camera simulator with reinforcement learning models. However, despite these efforts, the lack of synthetic datasets for detecting soccer players persists because current simulators and games are not specifically designed for computer vision research and therefore lack the functionality to generate labeled datasets.
In this paper, we release SoccerSynth-Detection, the first synthetic dataset for soccer player detection. We validated its efficacy using the Yolov8n model [21], comparing it with real-world datasets: Soccernet-Tracking [23] and SportsMoT [27]. Experimental results demonstrate that models trained on SoccerSynth-Detection achieve superior generalization compared to those trained on real datasets, particularly in handling motion blur. Pretraining with SoccerSynth-Detection further enhances model performance, especially in low-data scenarios. We have also released the dataset generator of SoccerSynth-Detection, providing a powerful tool for research requiring extensive data.
The contributions of this paper are as follows:
  1. 1.
    We release the SoccerSynth-Detection dataset, the first synthetic dataset specifically designed for soccer player detection.
  2. 2.
    We develop the simulator [39] to generate soccer detection datasets by enhancing camera features for detection and introducing random lighting and textures.
  3. 3.
    We demonstrate that SoccerSynth-Detection can replace and complement real-world datasets for training detection models.
2 Method
2.1 Simulator Construction
Due to the limitations of current football simulators (e.g., [38, 40, 41]) in supporting bounding box generation, we built upon the previous soccer stadium simulator [39] to set up a new center camera, with its configurations based on the positions, angles, and focal lengths found in SoccerNet-Tracking [23] and SportsMoT [27], as illustrated in Figure 1. Additionally, we incorporated dynamic spectator models purchased from the Unreal Engine Marketplace.
Figure 1: Soccer Stadium. From left to right: (a) Drone view (b) Corner camera view, and (c) Center camera view.
To enhance the realism of the synthetic players, we integrated character models and animations sourced from the Unreal Engine Marketplace and implemented their movement logic using the Behavior Tree and AI Controller within Unreal Engine. Additionally, we simulated actions such as passing, dribbling, shooting, and receiving through the physics engine, enabling continuous and autonomous football matches within the stadium environment.
The behavioral rules governing the players are as follows:
  • Possession Rules: Possession is assigned to the team of the last player who touched the football.
  • Actions in Possession: The goalkeeper passes to forwards; defenders to midfielders; midfielders to forwards. Forwards either pass back to midfielders if surrounded or dribble towards the penalty area and shoot at a randomly selected target within the goal.
  • Actions When Opponents Possess the Ball: The nearest player tries to regain possession. The goalkeeper acts to regain possession if the ball is within the penalty area. Other players engage in strategic off-ball movements.
  • Off-Ball Movement: Teams with possession move towards the opponent’s goal, while teams without possession retreat towards their own goal, excluding goalkeepers.
  • Reset Rules: Begin with 4-4-2 formation. The game includes no offside rule and resets when the football exits the field or the match time expires.
  • Football Movement: Simulated with a physics engine, where players kick the ball forward periodically as they dribble.
  • Passing Mechanics: Briefly stops the ball, calculates a parabolic trajectory, and applies the necessary force.
  • Receiving Mechanics: Momentarily stops the ball upon receipt to simulate control.
2.2 Data Generation
To enhance the diversity of the dataset, we first applied extensive randomization of lighting and textures during the rendering process, along with random camera rotations. Subsequently, we utilized the “2D Bounding Box for Deep Learning Image Detection” plugin purchased from the Unreal Engine Marketplace to capture frames and generate player detection bounding boxes. All individuals on the field were labeled as players, and the annotations were stored in YOLO format.
Models trained on clean and homogenous synthetic datasets often face significant performance drops when applied to real-world scenarios, primarily due to the domain gap between synthetic and real-world data. To address this issue, we drew on prior research on synthetic dataset generation and incorporate domain randomization techniques into the rendering process, as outlined in references [28, 29, 30, 31, 32, 33], to enhance the diversity and robustness of the generated data.
During the dataset generation process, we incorporated the following renderings:
  • Motion Blur Implementation: By introducing random camera shakes during filming, we restored the motion blur effect as illustrated in Figure 2a.
  • Field Characteristics Adjustment: In the simulated environment, we devised a combination of ten yellow and green grass colors, supplemented by five different grass textures, to increase the visual complexity of the background as illustrated in Figure 2b.
  • Lighting Conditions Simulation: To simulate varying times of day, we designed a gradient sunlight source system capable of adjusting brightness and color temperature, and emulating the rise and set of the sun as illustrated in Figure 2c.
  • Randomization of Player Apparel: By selecting randomly from a vast library of pre-made patterns, we achieved a random combination of hairstyles, jerseys, shorts, and shoes for players. This design aims to cover the diverse appearances of players in real-world scenarios.
  • Camera Setup Simulation: Camera angles and focal lengths were randomized, but optimized based on real-world game recordings to ensure the authenticity and practicality of viewing angles.
  • Player Actions and Scene Recreation: Utilizing behavior tree components, we simulated players’ movements such as running, shooting, and passing, recreating complex and dynamic match scenes.
Figure 2: Examples of SoccerSynth-Detection. From left to right: (a) Motion Blur (b) Field Characteristics (c) Lighting Conditions
Utilizing the “2D Bounding Box for Deep Learning Image Detection” plugin from the Unreal Engine Marketplace, we efficiently stored pixel-accurate detection bounding boxes for soccer balls and players in the captured frames during the rendering process, resulting in the creation of the dataset named SoccerSynth-Detection.
SoccerSynth-Detection contains two labeled classes, where 0 represents players and 1 represents soccer ball. It is divided into training and validation subsets and includes two types of images: clear images with a resolution of 1920×1080 and motion-blurred images with a resolution of 800×600. The bounding box annotations were saved in YOLO format.
2.3 Validation
To validate the effectiveness of our method, we generated 37,621 normal and 17,052 motion-blurred images using the simulator, dividing them into training and validation datasets at a 4:1 ratio, followed by conducting transfer learning and pretraining experiments. Both of the experiments were conducted with an image size of 640x640, a learning rate of 0.01, for up to 300 epochs, and an Intersection over Union (IoU) threshold of 0.7, with data augmentation applied. To ensure consistency with the actual annotations, we manually expanded the horizontal width of the detection boxes by two pixels. In both experiments, we utilized the YOLOv8n model [21] with the pretrained parameters “yolov8n.pt”, noted for its minimal parameter count. In both experiments, we evaluated the performance of the YOLOv8n model using Average Precision (AP), a metric that balances precision (the model’s ability to correctly identify relevant objects) and recall (the model’s ability to detect all relevant objects). A higher AP score reflects greater model accuracy. Specifically, AP50 measures AP at an IoU threshold of 0.5, providing insight into the model’s basic detection capabilities. In contrast, AP50:5:95 calculates the average AP across multiple IoU thresholds (ranging from 0.5 to 0.95 in increments of 0.05), offering a comprehensive evaluation of the model’s performance under varying detection strictness. In the fields of autonomous driving and pedestrian tracking, models trained on synthtic datasets have demonstrated superior generalization capabilities compared to those trained on real datasets [31, 28]. Therefore, we decided to conduct transfer experiments to evaluate the generalization performance of the model trained on SoccerSynth-Detection.
To facilitate this, we conducted a synth-to-real control experiment in which the YOLOv8n model was independently trained on the SportsMoT, SoccerNet-Tracking, and SoccerSynth-Detection datasets with all person-related labels standardized as “player”. Subsequently, the models underwent performance evaluations on previously unseen real-world datasets: SoccerNet-Tracking and SportsMoT, to ascertain their generalization capabilities in authentic scenarios. This experiment aimed to evaluate whether synthtic datasets can replace real datasets for training player detection algorithms in the field of soccer video analysis.
Limitations in the development of soccer video datasets restrict available data, impairing the effectiveness of player detection algorithms. Pre-training has proven to be a powerful method to improve algorithm performance, particularly in settings with limited data. Recent research initiatives increasingly focus on the application of synthtic datasets for pre-training purposes [42, 43]. Therefore, we decided to conduct pretraining experiments.
In the pretraining experiment, we investigated the potential of synthetic datasets to augment real-world datasets and enhance the accuracy of object detection models. Sub-datasets were created from the Soccernet-Tracking and SportsMoT datasets at different subsampling rates (1/2, 1/4, and 1/8) to evaluate the impact of training data volume on model performance. We compared the AP50 and AP50:5:95 metrics for the YOLOv8n model trained on the sub-datasets, both with and without the incorporation of the SoccerSynth-Detection dataset for pretraining.
3 Results
First, in the transfer experiment, we evaluated the generalization performance of the model trained on SoccerSynth-Detection. TAs illustrated in Table 1, while the Yolov8n model trained on the SoccerSynth-Detection dataset exhibited a minor decline of 2.2 percentage points at AP50, it demonstrated a gain of 1.0 percentage points at AP50:5:95. Furthermore, performance evaluations on the SoccerNet-Tracking dataset indicated substantially better outcomes. The Yolov8n model, when trained on the SoccerSynth-Detection dataset, achieved significant enhancements of 6.6 percentage points at AP50 and 4.5 percentage points at AP50:5:95.
Table 1: Performance comparison on the transfer experiment. The models were independently trained on the SportsMoT, SoccerNet-Tracking, and SoccerSynth-Detection datasets and tested on the SportsMoT and SoccerNet-Tracking datasets. The bold text indicates the best performance.
Training Data Test Data AP50 AP50:5:95
SoccerNet-Tracking SportsMoT 90.3 62.3
SoccerSynth-Detection SportsMoT 88.1 63.3
SportsMoT SoccerNet-Tracking 82.8 43.9
SoccerSynth-Detection SoccerNet-Tracking 89.4 48.4
[…]
Next, in the pretrainig experiment, we examined the performance in settings with limited data. TTable 2 demonstrates that when using SoccerNet-Tracking as the real-world dataset, the YOLOv8n model pretrained with SoccerSynth-Detection achieved an improvement of 0.1 to 1.5 percentage points in AP50 and AP50:5:95 metrics, effectively mitigating the adverse impact caused by reduced data volume. When SportsMOT was used as the real-world dataset with a subsampling rate of 1/8, the pretrained YOLOv8n model achieved the highest AP50 of 95.3 percent.
Table 2: Performance comparison in the pretraining experiment. Models were trained on SoccerNet-Tracking and SportsMoT Datasets with varying subsampling rates, evaluated both with and without pre-training on SoccerSynth-Detection. The bold text indicates the best performance
Training Data AP50 AP50:5:95
Dataset Subsampling Rate Without Pretrain With Pretrain Without Pretrain With Pretrain
SoccerNet-Tracking 1 96.8 97.0 62.2 62.8
1/2 96.8 96.9 61.8 62.4
1/4 96.4 96.9 61.1 61.9
1/8 96.2 96.5 60.5 62.0
SportsMoT 1 93.2 93.2 78.0 78.0
1/2 93.6 92.6 77.3 77.5
1/4 93.3 93.7 76.9 78.0
1/8 92.6 95.3 73.1 75.8
[…]
4 Discussion
Here we discuss the results of two experiments: transfer and pretraining experiments.
Figure 3: Transfer Experiment Results. From left to right: (a) Left: SoccerNet-Tracking, Right: SoccerSynth-Detection results with motion blur. (b) Top: SportsMoT, Bottom: SoccerSynth-Detection results with motion blur. (c) Left: SoccerNet-Tracking, Right: SoccerSynth-Detection performance degradation.
In the transfer experiment, we observed that the networks trained on the Soccersynth-Detection dataset outperformed those trained on other datasets in most cases, as demonstrated in Table 1. However, in comparisons where the test set was SportsMoT, although the Soccersynth-Detection outperformed SoccerNet-Tracking in the AP50:5:95 metric, it was inferior in the AP50 metric. Hence, we visualize the network’s predictions to analyze the underlying reasons for this discrepancy.
Figure 3a reveals that the Yolov8n network, when trained with SoccerNet-Tracking, produces multiple overlapping boxes with significant deviations on SportsMoT images under these conditions. In contrast, the same model trained on SoccerSynth displayed higher accuracy in identifying players. A similar trend was noted in tests with the SoccerNet-Tracking dataset, where the SportsMoT-trained Yolov8n failed in player detection, but the SoccerSynth counterpart excelled. The superior performance of the SoccerSynth-trained Yolov8n likely arose from its exposure to a greater variety and volume of motion blur images, enhancing its generalization across datasets. This ability to more precisely fit detection contours to player outlines also contributed to its outperformance in the AP50:5:95 metric relative to other networks. Figure 3c highlighted instances where the Yolov8n network, trained on SoccerSynth, fails to detect players’ heads. This issue likely arose from insufficient diversity in player head models within the SoccerSynth-Detection dataset, contributing to the lower AP50 performance on the SportsMoT dataset. In the future, we plan to introduce a more diverse array of head texture mappings to address this issue.
Figure 4: Results of experiments where the training dataset is SportsMoT and that with the subsampling rate of 1/8. Left: Result without pretraining, Right: Result with pretraining.
In the pretraining experiment, using SoccerSynth-Detection consistently enhanced model performance across different datasets and subsampling rates, particularly in settings with sparse training data as presented in Table 2. We displayed the prediction results of the network that showed the most improvement after pretraining, specifically when the training dataset was SportsMoT with a subsampling rate of 1/8, as illustrated in Figure 4. This figure highlighted that pretraining helped the model effectively avoid false detections and achieve higher confidence in detection boxes, offering a clear explanation for the observed overall performance boost post-pretraining.
Our results derived from pre-training and transfer learning suggest that the generalization capability of SoccerSynth-Detection is comparable to real-world datasets. It significantly enhances the performance of detection algorithms under certain conditions and proves to be a critical supplement to real datasets. This approach successfully assembles a broad array of annotated player detection videos at minimal cost, demonstrating its potential for wide application across various sports to address the shortage of video datasets with detection annotations. We believe this methodology will drive significant advancements in sports video analytics.
This study has two limitations: the first is the confinement to the central camera, and the second is the discontinuity between frames to increase data diversity. In future work, we plan to reference actual soccer match filming practices, incorporating cameras at various positions. Additionally, we will introduce distinctions among players and assign numbered labels to ensure frame continuity over certain periods.
5 Conclusion
In this paper, we addressed the current shortage of datasets for player detection by adding a dataset generation feature for player detection to our previous simulator, incorporating a large amount of random lighting, textures, and motion blur, and releasing the first synthetic player detection dataset, SoccerSynth-Detection. Through experiments, we demonstrated that SoccerSynth-Detection could replace real datasets or be used for pre-training to enhance the final performance of algorithms, especially outperforming existing datasets in scenarios with motion blur. Our work demonstrates the potential of synthetic datasets to replace real datasets for algorithm training in the field of soccer video analysis, which will help address the current shortage of athlete detection in sports video analytics.
Funding This work was financially supported by JST SPRING, Grant Number JPMJSP2125. The author H. Q. would like to take this opportunity to thank the “THERS Make New Standards Program for the Next Generation Researchers”.
Data availability statement The datasets generated and/or analyzed during the current study will be released after the publication.
Declarations
Conflict of Interest All authors declare that they have no conflict of interest.

### v2

@@
 Our work demonstrates the potential of synthetic datasets to replace real datasets for algorithm training in the field of soccer video analysis.
+The dataset is available at https://github.com/open-starlab/SoccerSynth-Detection.
 keywords:
@@
 First, in the transfer experiment, we evaluated the generalization performance of the model trained on SoccerSynth-Detection.
-TAs illustrated in Table 1, while the Yolov8n model trained on the SoccerSynth-Detection dataset exhibited a minor decline of 2.2 percentage points at AP50, it demonstrated a gain of 1.0 percentage points at AP50:5:95.
+As illustrated in Table 1, while the Yolov8n model trained on the SoccerSynth-Detection dataset exhibited a minor decline of 2.2 percentage points at AP50, it demonstrated a gain of 1.0 percentage points at AP50:5:95.
 Furthermore, performance evaluations on the SoccerNet-Tracking dataset indicated substantially better outcomes.
@@
 Next, in the pretrainig experiment, we examined the performance in settings with limited data.
-TTable 2 demonstrates that when using SoccerNet-Tracking as the real-world dataset, the YOLOv8n model pretrained with SoccerSynth-Detection achieved an improvement of 0.1 to 1.5 percentage points in AP50 and AP50:5:95 metrics, effectively mitigating the adverse impact caused by reduced data volume.
+Table 2 demonstrates that when using SoccerNet-Tracking as the real-world dataset, the YOLOv8n model pretrained with SoccerSynth-Detection achieved an improvement of 0.1 to 1.5 percentage points in AP50 and AP50:5:95 metrics, effectively mitigating the adverse impact caused by reduced data volume.
 When SportsMOT was used as the real-world dataset with a subsampling rate of 1/8, the pretrained YOLOv8n model achieved the highest AP50 of 95.3 percent.
@@
 Conflict of Interest All authors declare that they have no conflict of interest.
+Author Contribution H.Q. implement the system and H.Q. and K.F. wrote the main manuscript text. All authors discussed this research and reviewed the manuscript.